Each the interior angle of a regular polygon exceeds the exterior angle by 160 degree .find the number of sides of the polygon
Answer by stanbon(75887)

Each the interior angle of a regular polygon exceeds the exterior angle by 160 degree .find the number of sides of the polygon.
----
Let the exterior angle be "x".
The corresponding interior angle is "x+160"
Equation:
x + x+160 = 180
2x = 20
x = 10 degrees (measure of one of the exterior angles)
-----
Fact: The sum of all of the exterior angles is 360 degrees.
-----
# of exterior angles = 360/10 = 36
---
Therefore, # of sides is 36
=================================
Cheers,
Stan H.